OpenStudy (anonymous):

find the difference 6 5/8 - 1/6=

OpenStudy (anonymous):

The LCD is 24, you want to rewrite the fractions in terms of the LCD here: \[6\frac{5}{8}=6\frac{15}{24}\] 8*3=24, 5*3=15 \[\frac{1}{6}=\frac{4}{24}\] 6*4=24 1*4=4 Now you can subtract them easily

OpenStudy (anonymous):

\[6\frac{15}{24}-\frac{4}{24}=6\frac{15-4}{24}\]

OpenStudy (anonymous):

is it 65/8-1/6???

OpenStudy (anonymous):

6 11/24 ?

OpenStudy (anonymous):

Yes that's the right answer

OpenStudy (anonymous):

do the cross multiplication !!!!

OpenStudy (anonymous):